OpenStudy (anonymous):

rabbits have a weigth that is normally distributed with a mean of 80 pounds and standard deviation 13 pounds. Recent study claims average weight has increased. An animal is selected at random and its weight is 96 pounds. What is the p value

OpenStudy (kropot72):

Treating the calculation as a normal approximation to a discrete distribution, and applying the continuity correction, the z-scores are \[z _{1}=\frac{96-80-0.5}{13}=1.1923\] \[z _{2}=\frac{96-80+0.5}{13}=1.2692\] From a standard normal distribution table the cumulative probabilities corresponding to the above z-scores are \[p _{2}=0.8978\] \[p _{1}=0.8834\] \[p(96\ pounds)=p _{2}-p _{1}=0.8978-0.8834=\ ....\]